3. The monopolist with marginal cost of production of 40 sells to two distinct regions. In Region 1, demand is given by Q1=300-p1. In region2, it is given by Q2=180-p2.

1) Determine the optimal uniform price and output when discrimination is impossible.

2) Assume discrimination between the two regions is possible. What price will be set for each region? What quantity will be sold in each region?

3) How does the discriminatory price relate to the elasticity of demand in each country?
1
Expert's answer
2015-09-19T00:00:43-0400
3. The monopolist with marginal cost of production of 40 sells to two distinct regions. In Region 1, demand is given by Q1=300-p1. In region 2, it is given by Q2=180-p2.
1) Total demand is Qd = Q1 + Q2 = 480 - 2p.
The optimal uniform price will be in the point, where P = MC = 40, so the output Q =180 - 2*40 = 400 units, when discrimination is impossible.
2) If discrimination between the two regions is possible, then the higher price will be set for the region 1 and the quantity sold in this region will be lower.
3) The higher price has the lower elasticity of demand.